OpenStudy (anonymous):

Solve the following system of equations and show all work. y = x2 + 3 y = x + 5

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@dan815

OpenStudy (ipwnbunnies):

I think the easiest way is to equate the x part to each other (since y = y), bring all the terms to one side, and solve the quadratic equation.

OpenStudy (ipwnbunnies):

x + 5 = x^2 + 3

OpenStudy (anonymous):

ok that makes sense, but now what?

OpenStudy (ipwnbunnies):

Take all the terms to the right side, so the left side becomes 0.

OpenStudy (anonymous):

x+5-x^2-3=0 ?

OpenStudy (ipwnbunnies):

Yeah, like that. Although, it might've been easier to put the left part on the right side, so you won't have a negative x^2 term. But it makes no difference. If you multiply both sides by -1, you'll get the nicer form. Do that, and simplify.

OpenStudy (anonymous):

wait I'm confused, can you show me?

OpenStudy (anonymous):

oh x^2+3-x+5 =0 ???

OpenStudy (ipwnbunnies):

0 = x^2 + 3 - x - 5

OpenStudy (ipwnbunnies):

X and 5 are both positive on the left side, so when you subtract it, it'll be negative on the right side.

OpenStudy (anonymous):

ok got it, no combine 3 and -5 ?

OpenStudy (anonymous):

*now

OpenStudy (ipwnbunnies):

Exactly! Finally, solve the quadratic equation.

OpenStudy (anonymous):

available to help with one more?

OpenStudy (ipwnbunnies):

I can't find now unfortunately. I'm sorry :c But good luck!

OpenStudy (ipwnbunnies):

right now*

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait can you show me the equation real quick with 3 and -5 combined?

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

nevermind sorry